OpenStudy (anonymous):

Find the slope of a line that is parallel to the given line,if the slope is undefined,state this y=-3

OpenStudy (anonymous):

What is the given line?

OpenStudy (anonymous):

Is it y=-3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

That's jsut a straight horizontal line

OpenStudy (anonymous):

y=any number would be parallel

OpenStudy (anonymous):

Are you sure you wrote the question out correctly?

OpenStudy (anonymous):

I solved it and this is what I got the slope of the line is -3 and on a horinontal line -3 will end up been Y=3

OpenStudy (anonymous):

y=-3 is not a slope

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Yes its correct here the question again Find the slope of a line that is parallel to the given line,if the slope is undefined,state this y=-3

OpenStudy (anonymous):

They want me to state it if the slope is undefined

OpenStudy (anonymous):

b=-3 in that scenario

OpenStudy (anonymous):

and slope is 0

OpenStudy (anonymous):

The slope is defined, It is 0. No slope is still a slope.

OpenStudy (anonymous):

y=-3 is y=(0)x-3

OpenStudy (anonymous):

since y=mx+b is equation of a line, in that case, m=0 and m is slope

OpenStudy (anonymous):

Oh good I understand it

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

So the slope of the line that is parallel to y=-3 is (0)x-3

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

The EQUATION of the line is y=0x-3. The actual slope is 0. The slope is part of the equation. Remember y=mx+b is the equation of a line. The "m" is the slope. So the actual slope, is m=0. The "b" is vertical translation up and down and since it's y=-3, it's a vertical translation down 3 units. So slope = m = 0.

OpenStudy (anonymous):

And y=-3 is a straight line crossing y at -3 with no slope upwards or downwards.

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Great .

OpenStudy (anonymous):

U sure u understand?

OpenStudy (anonymous):

I have 1 more problem : Find the slope of a line that is parallel to the given line,if the slope is undefined,state this y=-1/5x+2

OpenStudy (anonymous):

Post in new question